# Convergence, Duality and Well-Posedness in Convex Bilevel Optimization

### v1

@@ 3 Conditions for strong duality and convergence
We now provide two conditions to ensure that strong duality holds for Eq. 2, thereby (equivalently) ensuring that any algorithm which guarantees Eq. 5 has assurance that Eq. 7 holds also. {condition} If $\{x_{t}\}_{t\geq 1}\subset X$ is a sequence such that $g(x_{t})\to g^{*}$, we have $\operatorname{Dist}(x_{t},X_{g})\to 0$.
{condition}
@@
Section 3 generalizes, for example, error bound-type conditions on $g$, where $\tau\operatorname{Dist}(x,X_{g})^{r}\leq g(x)-g^{*}$. Error bounds are used in several works [e.g., 16, 5, 20] to enable stronger guarantees for algorithms solving Eq. 2.
Section 3 can be shown to be equivalent to the coercivity of the functions $h=\max\{f,g\}$ and $\bar{h}=\max\{f-f^{*},g-g^{*}\}$ over $X$. (Recall that a function $h$ is coercive over $X$ if $h(x)\to\infty$ along any recessive direction in $X$.) Section 3, or more restrictive versions of it, has appeared in various works. Cabot [3], Solodov [26, 27], Malitsky [18], Dempe et al. [9] assume exactly Section 3. Helou and Simões [14], Amini and Yousefian [1], Kaushik and Yousefian [17], Shen et al. [24], Giang-Tran et al. [13], Jiang et al. [16], Cao et al. [5] employ the assumption that $X$ is bounded to prove guarantees for their algorithms. This immediately implies Section 3 holds. Finally, Beck and Sabach [2], Sabach and Shtern [21], Shehu et al. [23], Doron and Shtern [10], Merchav and Sabach [19] use the assumption that $f$ itself is coercive, which also implies Section 3. The coercivity assumption on $f$ is commonly motivated by learning problems where $g$ models the prediction error and $f$ is a regularization function, e.g., the $\ell_{1}$- or $\ell_{2}$-norm.
Section 3 guarantees that strong duality holds, and hence Eq. 5 implies Eq. 7.
Proposition 3.1.
Suppose that Section 3 holds for Eq. 2. Then strong duality holds, and, consequently, any sequence satisfying Eq. 5 also satisfies Eq. 7.
Proof 3.2 (Proof of Proposition 3.1).
Suppose that $\{x_{t}\}_{t\geq 1}$ is a sequence which satisfies the premise of Lemma 2.4: $g(x_{t})\to g^{*}$ and $\limsup_{t\to\infty}f(x_{t})\leq d^{*}$. Let $x^{*}\in X^{*}$ be an optimal solution to Eq. 2, and define the sequence $\{y_{t}\}_{t\geq 1}\subset X_{g}$ by $y_{t}:=\operatorname*{arg\,min}_{y\in X_{g}}\|x_{t}-y\|_{2}$. (Note that the minimizer is unique since the $\ell_{2}$-norm is strictly convex.) By convexity of $f$, we have
@@
The second inequality follows from the Cauchy-Schwarz inequality, while the third follows since $\nabla f(x^{*})^{\top}(y_{t}-x^{*})\geq 0$ by the optimality of $x^{*}$ in minimizing $f$ over $X_{g}$, and the fact that $y_{t}\in X_{g}$. By Section 3, we have $\|x_{t}-y_{t}\|_{2}=\operatorname{Dist}(x_{t},X_{g})\to 0$ since $g(x_{t})\to g^{*}$. Therefore, the limit inferior of both sides is
@@
so strong duality holds.
Unfortunately, Section 3 on its own is not enough to guarantee that Eq. 8 holds.
Example 3.3.
Define $X=\{(x_{1},x_{2},x_{3}):x_{1}\geq 1,0\leq x_{2}\leq 1\}\subset{\mathbb{R}}^{3}$, and $g(x)=x_{3}^{2}$. Then $g^{*}=0$ and $X_{g}=\{(x_{1},x_{2},x_{3}):x_{1}\geq 1,0\leq x_{2}\leq 1,x_{3}=0\}$. Notice that for any $x\in X$, $\operatorname{Dist}(x,X_{g})=|x_{3}|=\sqrt{g(x)}$, therefore Section 3 is satisfied. Now define $f(x)=x_{2}^{2}/x_{1}$, which is convex. Then $f^{*}=0$ and $X^{*}=\{(x_{1},x_{2},x_{3}):x_{1}\geq 1,x_{2}=0,x_{3}=0\}$. Now consider the sequence $\{x_{t}\}_{t\geq 1}$ defined by $x_{t}=(t,1,0)\in X_{g}$. Then $f(x_{t})=1/t\to f^{*}$, $g(x_{t})=0=g^{*}$, $\operatorname{Dist}(x_{t},X_{g})=0$ but $\operatorname{Dist}(x_{t},X^{*})=1$ for all $t$. Therefore, Eq. 8 is not satisfied.
Section 3 is more powerful than Section 3, and in fact has strong connections to the study of well-posedness, which examines properties of approximation schemes for solutions of optimization (and related) problems.
Definition 3.4 (Huang and Yang [15]).
@@
  • Levitin-Polyak (LP) well-posedness if, for any sequence $\{x_{t}\}_{t\geq 1}\subset X$ which satisfies $\operatorname{Dist}(x_{t},X_{g})\to 0$ and $f(x_{t})\to f^{*}$, there exists a subsequence $\{x_{t_{k}}\}_{k\geq 1}$ such that $x_{t_{k}}\to\bar{x}$ for some $\bar{x}\in X^{*}$.
  • generalized LP well-posedness if, for any sequence that satisfies Eq. 7, there exists a subsequence $\{x_{t_{k}}\}_{k\geq 1}$ such that $x_{t_{k}}\to\bar{x}$ for some $\bar{x}\in X^{*}$.
  • strong generalized LP well-posedness if, for any sequence that satisfies Eq. 5, there exists a subsequence $\{x_{t_{k}}\}_{k\geq 1}$ such that $x_{t_{k}}\to\bar{x}$ for some $\bar{x}\in X^{*}$.
It is immediately clear that strong generalized LP well-posedness implies generalized LP well-posedness. We will show that every sense of well-posedness in Definition 3.4 implies Section 3. It turns out that strong generalized LP well-posedness is equivalent to Section 3 in finite dimensions (which is our setting), which follows from a result of Huang and Yang [15]. We will additionally show that Section 3 ensures that strong duality holds, and thus Theorem 2.3 implies that generalized LP is equivalent to strong generalized LP well-posedness for Eq. 2.
Proposition 3.5.
@@
    Eq. 2 is strong generalized LP well-posed.
Furthermore, if Eq. 2 is LP well-posed, then Section 3 holds.
Proof 3.6 (Proof of Proposition 3.5).
First, note that if $X^{*}$ is not compact, then it must be unbounded, since we assume that $X^{*}$ is closed. Take a sequence $\{x_{t}\}_{t\geq 1}\subset X^{*}$ such that $\|x_{t}\|_{2}\to\infty$. Then this sequence satisfies all three premises of Definition 3.4, however there is no convergent subsequence. Therefore all three senses of well-posedness imply compactness of $X^{*}$, i.e., Section 3. Huang and Yang [15, Theorem 2.4] states that when $X$ is finite-dimensional, which we assume, then strong generalized LP well-posedness is equivalent to Section 3.
To show that strong generalized LP is equivalent to generalized LP, we will first show that under Section 3, strong duality holds. First, denote
$X_{1}:=\left\{x\in X:~f(x)-f^{*}\leq 1,~g(x)-g^{*}\leq 1\right\}.$
Note that $X_{1}$ is the $1$-sublevel set of the function $\bar{h}=\max\{f-f^{*},g-g^{*}\}$, and $X^{*}$ is the 0-sublevel set. Since $X^{*}$ is bounded, all sublevel sets of $\bar{h}$ are bounded, hence $X_{1}$ is also bounded. Let $\{x_{t}\}_{t\geq 1}\subset X$ be a sequence satisfying the premise of Lemma 2.4, i.e., $g(x_{t})\to g^{*}$ and $\limsup_{t\to\infty}f(x_{t})\leq d^{*}$. By weak duality $d^{*}\leq f^{*}$, therefore $f(x_{t})\leq f^{*}+1$ eventually, and also $g(x_{t})\leq g^{*}+1$ eventually as well. This means that only finitely many $x_{t}\not\in X_{1}$, and since $X_{1}$ is bounded, the whole sequence is bounded. Take a convergent subsequence $\{x_{t_{k}}\}_{k\geq 1}$ such that $x_{t_{k}}\to\bar{x}$. Since $X$ is closed, $\bar{x}\in X$. Furthermore, since $g$ is convex, it is lower semicontinuous, so $g^{*}=\liminf_{k\to\infty}g(x_{t_{k}})\geq g(\bar{x})\geq g^{*}$, hence $\bar{x}\in X_{g}$. But this implies that $f(\bar{x})\geq f^{*}$ by definition of $f^{*}$. Therefore $f^{*}\leq\limsup_{t\to\infty}f(x_{t})\leq d^{*}$ (where the last inequality follows from the premise on the sequence $\{x_{t}\}_{t\geq 1}$), and thus strong duality holds.
Now, assume that Eq. 2 is generalized LP well-posed. Then $X^{*}$ must be compact. Take a sequence $\{x_{t}\}_{t\geq 1}\subset X$ satisfying Eq. 5. Since compactness of $X^{*}$ implies strong duality, from Theorem 2.3 we know that $\{x_{t}\}_{t\geq 1}$ also satisfies Eq. 7. Thus, by generalized LP well-posedness, there exists $\bar{x}\in X^{*}$ and a subsequence for which $x_{t_{k}}\to\bar{x}$, hence Eq. 2 is also strong generalized LP well-posed. Since a strong generalized LP well-posed problem is also generalized LP well-posed, this shows the equivalence.
This equivalence yields a quantitative version of Eq. 8.
Corollary 3.7.
Suppose that Section 3 holds for Eq. 2. Then there exists a function $c:{\mathbb{R}}_{+}^{2}\to{\mathbb{R}}_{+}$ such that $c(0,0)=0$ and for any $x\in X$,
@@
Furthermore, the function $c$ satisfies that whenever $\beta_{t}\to 0$, $c(\alpha_{t},\beta_{t})\to 0$, then $\alpha_{t}\to 0$. Therefore, under Section 3 any sequence $\{x_{t}\}_{t\geq 1}$ satisfying Eq. 5 also satisfies Eq. 8.
Proof 3.8 (Proof of Corollary 3.7).
Proposition 3.5 states that Section 3 is equivalent to generalized LP well-posedness. Huang and Yang [15, Theorem 2.2] then ensures that whenever Eq. 2 is generalized LP well-posed, the function $c$ that satisfies the required properties is
$c(\alpha,\beta):=\inf_{x}\left\{|f(x)-f^{*}|:~x\in X,~\operatorname{Dist}(x,X^{*})=\alpha,~g(x)-g^{*}=\beta\right\}.$

### v2

@@ 3 Conditions for strong duality and convergence
We now provide two conditions to ensure that strong duality holds for Eq. 2, thereby (equivalently) ensuring that any algorithm which guarantees Eq. 5 has assurance that Eq. 7 holds also.
Condition 1.
If $\{x_{t}\}_{t\geq 1}\subset X$ is a sequence such that $g(x_{t})\to g^{*}$, we have $\operatorname{Dist}(x_{t},X_{g})\to 0$.
Condition 2.
@@
1 generalizes, for example, error bound-type conditions on $g$, where $\tau\operatorname{Dist}(x,X_{g})^{r}\leq g(x)-g^{*}$. Error bounds are used in several works [e.g., 16, 5, 20] to enable stronger guarantees for algorithms solving Eq. 2.
2 can be shown to be equivalent to coercivity of the functions $h=\max\{f,g\}$ and $\bar{h}=\max\{f-f^{*},g-g^{*}\}$. 2, or more restrictive versions of it, has appeared in various works. Cabot [3], Solodov [26, 27], Malitsky [18], Dempe et al. [9] assume exactly 2. Helou and Simões [14], Amini and Yousefian [1], Kaushik and Yousefian [17], Shen et al. [24], Giang-Tran et al. [13], Jiang et al. [16], Cao et al. [5] employ the assumption that $X$ is bounded to prove guarantees for their algorithms. This immediately implies 2 holds. Finally, Beck and Sabach [2], Sabach and Shtern [21], Shehu et al. [23], Doron and Shtern [10], Merchav and Sabach [19] use the assumption that $f$ is coercive, which also implies 2. This is commonly motivated by learning problems where $g$ models the prediction error and $f$ is a regularization function, e.g., the $\ell_{1}$- or $\ell_{2}$-norm.
1 guarantees that strong duality holds, and hence Eq. 5 implies Eq. 7.
Proposition 1.
Suppose that 1 holds for Eq. 2. Then strong duality holds, and, consequently, any sequence satisfying Eq. 5 also satisfies Eq. 7.
Proof of Proposition 1.
Suppose that $\{x_{t}\}_{t\geq 1}$ is a sequence satisfying the premise of Lemma 1: $g(x_{t})\to g^{*}$ and $\limsup_{t\to\infty}f(x_{t})\leq d^{*}$. Let $x^{*}\in X^{*}$ be an optimal solution to Eq. 2, and define the sequence $\{y_{t}\}_{t\geq 1}\subset X_{g}$ by $y_{t}:=\operatorname*{arg\,min}_{y\in X_{g}}\|x_{t}-y\|_{2}$. (Note that the minimizer is unique since the $\ell_{2}$-norm is strictly convex.) By convexity of $f$, we have
@@
The second inequality follows from the Cauchy-Schwarz inequality, while the third follows since $\nabla f(x^{*})^{\top}(y_{t}-x^{*})\geq 0$ by the optimality of $x^{*}$ in minimizing $f$ over $X_{g}$, and the fact that $y_{t}\in X_{g}$. By 1, we have $\|x_{t}-y_{t}\|_{2}=\operatorname{Dist}(x_{t},X_{g})\to 0$ since $g(x_{t})\to g^{*}$. Therefore, the limit inferior of both sides is
@@
so strong duality holds. ∎
Unfortunately, 1 on its own is not enough to guarantee that Eq. 8 holds.
Example 2.
Define $X=\{(x_{1},x_{2},x_{3}):x_{1}\geq 1,0\leq x_{2}\leq 1\}\subset{\mathbb{R}}^{3}$, and $g(x)=x_{3}^{2}$. Then $g^{*}=0$ and $X_{g}=\{(x_{1},x_{2},x_{3}):x_{1}\geq 1,0\leq x_{2}\leq 1,x_{3}=0\}$. Notice that for any $x\in X$, $\operatorname{Dist}(x,X_{g})=|x_{3}|=\sqrt{g(x)}$, therefore 1 is satisfied. Now define $f(x)=x_{2}^{2}/x_{1}$, which is convex. Then $f^{*}=0$ and $X^{*}=\{(x_{1},x_{2},x_{3}):x_{1}\geq 1,x_{2}=0,x_{3}=0\}$. Now consider the sequence $\{x_{t}\}_{t\geq 1}$ defined by $x_{t}=(t,1,0)\in X_{g}$. Then $f(x_{t})=1/t\to f^{*}$, $g(x_{t})=0=g^{*}$, $\operatorname{Dist}(x_{t},X_{g})=0$ but $\operatorname{Dist}(x_{t},X^{*})=1$ for all $t$. Therefore Eq. 8 is not satisfied.
2 is more powerful than 1, and in fact has strong connections to the study of well-posedness, which examines properties of approximation schemes for solutions of optimization (and related) problems.
Definition 2 (Huang and Yang [15]).
@@
  • Levitin-Polyak (LP) well-posedness if, for any sequence $\{x_{t}\}_{t\geq 1}\subset X$ which satisfies $\operatorname{Dist}(x,X_{g})\to 0$ and $f(x_{t})\to f^{*}$, there exists a subsequence $\{x_{t_{k}}\}_{k\geq 1}\subset\{x_{t}\}_{t\geq 1}$ and $\bar{x}\in X^{*}$ such that $x_{t_{k}}\to\bar{x}$.
  • generalized LP well-posedness if, for any sequence that satisfies Eq. 7, there exists a subsequence $\{x_{t_{k}}\}_{k\geq 1}\subset\{x_{t}\}_{t\geq 1}$ and $\bar{x}\in X^{*}$ such that $x_{t_{k}}\to\bar{x}$.
  • strong generalized LP well-posedness if, for any sequence that satisfies Eq. 5, there exists a subsequence $\{x_{t_{k}}\}_{k\geq 1}\subset\{x_{t}\}_{t\geq 1}$ and $\bar{x}\in X^{*}$ such that $x_{t_{k}}\to\bar{x}$.
It is immediately clear that strong generalized LP well-posedness implies generalized LP well-posedness. We will show that every sense of well-posedness in Definition 2 implies 2. It turns out that strong generalized LP well-posedness is equivalent to 2 in finite dimensions (which is our setting), which follows from a result of Huang and Yang [15]. We will additionally show that 2 ensures that strong duality holds, and thus Theorem 1 implies that generalized LP is equivalent to strong generalized LP well-posedness for Eq. 2.
Proposition 2.
@@
    Eq. 2 is strongly generalized LP well-posed.
Furthermore, if Eq. 2 is LP well-posed, then 2 holds.
Proof of Proposition 2.
First, note that if $X^{*}$ is not compact, then it must be unbounded, since we assume that $X^{*}$ is closed. Take a sequence $\{x_{t}\}_{t\geq 1}\subset X^{*}$ such that $\|x_{t}\|_{2}\to\infty$. Then this sequence satisfies all three premises of Definition 2, however there is no convergent subsequence. Therefore all three senses of well-posedness imply compactness of $X^{*}$, i.e., 2. Huang and Yang [15, Theorem 2.4] states that when $X$ is finite-dimensional, which we assume, then strong generalized LP well-posedness is equivalent to 2.
To show that strong generalized LP is equivalent to generalized LP, we will first show that under 2, strong duality holds. First, denote
$X_{1}:=\left\{x\in X:f(x)-f^{*}\leq 1,g(x)-g^{*}\leq 1\right\}.$
Note that $X_{1}$ is the $1$-sublevel set of the function $\bar{h}=\max\{f-f^{*},g-g^{*}\}$, and $X^{*}$ is the 0-sublevel set. Since $X^{*}$ is bounded, all sublevel sets of $\bar{h}$ are bounded, hence $X_{1}$ is also bounded. Let $\{x_{t}\}_{t\geq 1}\subset X$ be a sequence satisfying the premise of Lemma 1, i.e., $g(x_{t})\to g^{*}$ and $\limsup_{t\to\infty}f(x_{t})\leq d^{*}$. By weak duality $d^{*}\leq f^{*}$, therefore $f(x_{t})\leq f^{*}+1$ eventually, and also $g(x_{t})\leq g^{*}+1$ eventually as well. This means that only finitely many $x_{t}\not\in X_{1}$, and since $X_{1}$ is bounded, the whole sequence is bounded. Take a convergent subsequence $\{x_{t_{k}}\}_{k\geq 1}$ such that $x_{t_{k}}\to\bar{x}$. Since $X$ is closed, $\bar{x}\in X$. Furthermore, since $g$ is convex, it is lower semicontinuous, so $g^{*}=\liminf_{k\to\infty}g(x_{t_{k}})\geq g(\bar{x})\geq g^{*}$, hence $\bar{x}\in X_{g}$. But this implies that $f(\bar{x})\geq f^{*}$ by definition of $f^{*}$. Therefore $f^{*}\leq\limsup_{t\to\infty}f(x_{t})\leq d^{*}$, and thus strong duality holds.
Now, assume that Eq. 2 is generalized LP well-posed. Then $X^{*}$ must be compact. Take a sequence $\{x_{t}\}_{t\geq 1}\subset X$ satisfying Eq. 5. Since compactness of $X^{*}$ implies strong duality, Theorem 1, we know that $\{x_{t}\}_{t\geq 1}$ also satisfies Eq. 7. Therefore by generalized LP well-posedness, there exists $\bar{x}\in X^{*}$ and a subsequence for which $x_{t_{k}}\to\bar{x}$, hence Eq. 2 is also strongly generalized LP well-posed. Since a strongly generalized LP well-posed problem is also generalized LP well-posed, this shows the equivalence. ∎
We now use this equivalence to provide a quantitative version of Eq. 8.
Corollary 1.
Suppose that 2 holds for Eq. 2. Then there exists a function $c:{\mathbb{R}}_{+}^{2}\to{\mathbb{R}}_{+}$ such that $c(0,0)=0$ and for any $x\in X$,
@@
Furthermore, the function $c$ satisfies that whenever $\beta_{t}\to 0$, $c(\alpha_{t},\beta_{t})\to 0$, then $\alpha_{t}\to 0$. Therefore, under 2 any sequence $\{x_{t}\}_{t\geq 1}$ satisfying Eq. 5 also satisfies Eq. 8.
Proof of Corollary 1.
Proposition 2 states that 2 is equivalent to generalized LP well-posedness. Huang and Yang [15, Theorem 2.2] then ensures that whenever Eq. 2 is generalized LP well-posed, the function $c$ that satisfies the required properties is
$c(\alpha,\beta):=\inf_{x}\left\{|f(x)-f^{*}|:x\in X,\operatorname{Dist}(x,X^{*})=\alpha,g(x)-g^{*}=\beta\right\}.$
4 Conclusion
We studied convergence guarantees for the convex bilevel problem (2). Through Example 1 we showed that typical bounds (5) are insufficient: they can hold while neither the objective values nor the iterates approach a bilevel solution in a meaningful sense. Our main result establishes an exact remedy: strong duality for the value-function formulation (3) is equivalent to the property that every sequence satisfying (5) necessarily enjoys objective convergence (7) (Theorem 1).
We then identified two simple conditions ensuring this regularity. First, the proximity assumption 1 guarantees strong duality and hence (7); however, as Example 2 demonstrates, it does not ensure convergence to the optimal solution set (8). Second, the compactness assumption 2 is stronger but widely utilized in previous work. It yields strong duality of Eq. 2 and solution-set convergence (Proposition 2) together with quantifiable bounds. These findings provide algorithm-agnostic criteria under which iterate sequences for Eq. 2 have meaningful limits.
Interestingly, Proposition 2 and Example 2 shows that Eq. 2 may still be ill-posed (in the sense of Definition 2) even when strong duality holds. Future work includes investigating the implications of the stronger notion of dual solvability on approximating sequences for Eq. 2. Furthermore, obtaining exact forms of the quantified bound in Corollary 1 for important bilevel problems is of interest.